OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, –1) with slope 4. A. y = 4x – 1 B. y = 4x – 25 C. y + 1 = 4(x – 6) D. y + 6 = 4(x – 1)

Directrix (directrix):

The general equation for the slope-intercept of a line is y = mx + b where b is the y-intercept and m is the slope of the line.

Directrix (directrix):

In this problem, m = 4 and the line passes through (6.-1). y = mx + b y = 4x + b

Directrix (directrix):

y = 4x + b From the point coordinates, x = 6 and y = -1 So, (-1) = 4*(6) + b

Directrix (directrix):

(-1) = 4*(6) + b Solve the above equation for b. --> @Melody_

OpenStudy (anonymous):

thankssss this time i completely understand this!!!@Directrix

OpenStudy (anonymous):

:D

Directrix (directrix):

Alrighty, then.